# I4VGen: Image as Stepping Stone for Text-to-Video Generation

### v1

@@ 4.2 Anchor image-guided video synthesis
  • (a) Rather than resampling the Gaussian noise at each iteration as in traditional SDS, we maintain a constant noise across the optimization, enhancing convergence speed.
  • (b) Optimization is confined to the initial stages of the denoising process, where noise levels are higher, focusing on dynamic information distillation.
  • (c) We implement a coarse-to-fine optimization strategy, evolving from high to low noise levels, specifically from timestep $T$ to $\tau_{\text{NI-VSDS}}$, where $T>\tau_{\text{NI-VSDS}}>0$. This approach stabilizes the optimization trajectory and yields superior motion quality.
The algorithmic implementation of our Noise-Invariant Video Score Distillation Sampling is detailed in Algorithm 1, which outlines the procedure for transitioning a static video into a dynamic one using the defined NI-VSDS loss. Remarkably, we only perform a single update from timestep $T$ to $\tau_{\text{NI-VSDS}}$, which requires fewer than 50 iterations, a significant reduction compared to the thousands typically necessary in SDS for text-to-3D synthesis. $\alpha$ is a scalar defining the step size of the gradient update. We empirically set $\tau_{\text{NI-VSDS}}=\mathsf{Int}(T\times p_{\text{NI-VSDS}})$.
Input: A T2V diffusion model $\epsilon_{\theta}(\cdot)$, a text prompt $\mathbf{y}$, a static video $\hat{\mathbf{v}}$, timestep $\tau_{\text{NI-VSDS}}$.
Output: A dynamic video.
@@ 3       $\hat{\mathbf{v}}_{t}\leftarrow\mathsf{AddNoise}\left(\hat{\mathbf{v}},\epsilon,t\right)$
Algorithm 1 Noise-invariant Video SDS
Video regeneration. Upon animating the static video, we further enhance the video quality through a diffusion-denoising process. This phase does not suffer from the aforementioned training-inference gap, thus enabling more refined generation outcomes.
Notably, we have the flexibility to introduce noise at any timestep $\tau_{\text{re}}$, calculated as $\tau_{\text{re}}=\mathsf{Int}(T\times p_{\text{re}})$, followed by the corresponding denoising process. This strategy not only preserves the fine appearance textures but also reduces the required denoising steps, thus streamlining the video synthesis process and elevating the overall quality of the resulting video.
@@ 5.1 Experimental settings
Implementation details. I4VGen is a training-free and plug-and-play strategy that can be adapted to existing text-to-video diffusion models. To ascertain the efficacy and adaptability of I4VGen, it iså applied to two well-regarded diffusion-based text-to-video models: AnimateDiff [25] and LaVie [73].
For AnimateDiff, the mm-sd-v15-v2 motion module [^1] https://github.com/guoyww/AnimateDiff, alongside Stable Diffusion v1.5, is utilized to synthesize 16 consecutive frames at a resolution of 512$\times$512 pixels for evaluation. For LaVie, the base-version [^2] https://github.com/Vchitect/LaVie is employed to generate 16 consecutive frames at 320$\times$512 pixels for evaluation. All other inference details adhere to the original settings described in [25] and [73], respectively. Notably, both AnimateDiff and LaVie possess inherent text-to-image capabilities when excluding the motion module. To avoid introducing additional GPU storage requirements, we leverage their corresponding image versions for text-to-image generation in I4VGen. For AnimteDiff, we empirically set $N$ = 16, $p_{\text{NI-VSDS}}$ = 0.4, $\alpha$ = 1, and $p_{\text{re}}$ = 1. For LaVie, we empirically set $N$ = 16, $p_{\text{NI-VSDS}}$ = 0.4, $\alpha$ = 1, and $p_{\text{re}}$ = 0.8. All experiments are conducted on a single NVIDIA V100 GPU (32 GB).
Benchmark. I4VGen is assessed using VBench [35], a comprehensive benchmark that evaluates video generation models across 16 disentangled dimensions, which is more authoritative than FVD. These dimensions provide a detailed analysis of generation quality from two overarching perspectives: Video Quality [^3] Video Quality includes 7 evaluation dimensions: Subject Consistency, Background Consistency, Temporal Flickering, Motion Smoothness, Dynamic Degree, Aesthetic Quality, and Imaging Quality. The first 5 evaluate temporal quality, and the last 2 evaluate frame-wise quality., focusing on the perceptual quality of the generated videos, and Video-Condition Consistency [^4] Video-Condition Consistency includes 9 evaluation dimensions: Object Class, Multiple Objects, Human Action, Color, Spatial Relationship, Scene, Appearance Style, Temporal Style, Overall Consistency. The first 6 evaluate semantics, the 7 and 8-th evaluate style, and the 9-th evaluates overall consistency., assessing how well the generated videos align with the provided conditions.
AnimateDiff [25] 87.11 95.22 95.99 93.12 74.89 56.07 64.29 83.69
AnimateDiff + FreeInit [77] 90.45 96.57 96.89 95.66 70.17 59.25 63.51 87.55
AnimateDiff + I4VGen 95.17 97.73 98.51 96.45 57.72 64.68 66.18 92.59
LaVie [73] 91.65 96.30 98.03 95.73 71.94 59.64 65.13 91.25
LaVie + FreeInit [77] 92.32 96.35 98.06 95.83 71.11 59.41 63.89 89.13
LaVie + I4VGen 94.12 96.90 98.55 96.37 70.55 60.88 66.55 92.26
AnimateDiff [25] 22.61 90.40 81.73 31.55 45.61 24.40 24.49 25.71
AnimateDiff + FreeInit [77] 26.92 93.00 86.39 30.71 44.61 23.98 25.03 25.61
AnimateDiff + I4VGen 57.22 95.80 91.98 45.20 54.67 25.07 26.11 28.01
LaVie [73] 24.02 94.80 83.64 26.27 52.89 23.67 24.94 27.25
LaVie + FreeInit [77] 22.59 94.20 84.34 27.46 52.70 23.61 24.85 26.89
LaVie + I4VGen 32.77 96.20 88.59 33.81 55.64 24.35 25.62 27.68
5.2 Qualitative comparison
Fig. 4 presents a comparative analysis of our results against state-of-the-art counterparts using identical text prompts and random seeds. I4VGen excels in enhancing both the temporal consistency and the frame-wise quality, alongside superior alignment with the text prompts. For instance, in the case of playing guitar, AnimateDiff suffers from poor video quality, and FreeInit encounters an incomplete guitar in the middle of the video. In contrast, our method effectively addresses these issues, maintaining stable temporal consistency. Furthermore, while baseline methods struggle with accurate synthesis of all text-described components, e.g., NYC Times Square and suit, I4VGen generates videos that are visually realistic and closely aligned with the text prompts by using intermediate images with a generation-selection mechanism.
@@ 5.2 Qualitative comparison
w/o generation-selection 94.89 97.80 98.28 96.99 55.91 62.23 64.18 90.95
w/o NI-VSDS 96.47 98.82 98.99 97.56 28.24 65.17 65.52 92.66
AnimateDiff + I4VGen 95.17 97.73 98.51 96.45 57.72 64.68 66.18 92.59
@@
w/o generation-selection 40.68 94.40 90.55 37.79 53.72 24.76 26.03 26.62
w/o NI-VSDS 62.84 94.80 91.95 47.57 55.80 24.88 25.72 27.91
AnimateDiff + I4VGen 57.22 95.80 91.98 45.20 54.67 25.07 26.11 28.01
@@ 5.3 Quantitative comparison
Objective evaluation. Following the protocols established by VBench [35], we evaluate I4VGen in terms of both video quality and video-text consistency. As detailed in Table 1, I4VGen outperforms all other approaches in temporal quality (higher background and subject consistency, less flickering, and better smoothness), frame-wise quality (higher aesthetic and imaging quality), and video-text consistency (greater semantics, style, and overall consistency). Although counterparts occasionally produce videos with more dynamic motion, they are often linked to inappropriate or excessive movements, as evidenced by the unnatural motion of a rabbit portrait shown in Fig. 4. I4VGen strikes a more effective balance between motion intensity and overall video quality, which is further verified in the user study.
Table 3: User study. I4VGen performs over other counterparts.
Methods User study
AnimateDiff [25] 6.67%
AnimateDiff + FreeInit [77] 22.67%
AnimateDiff + I4VGen 70.67%
LaVie [73] 24.00%
LaVie + FreeInit [77] 17.33%
LaVie + I4VGen 58.67%
User study. We conduct a subjective user study involving 10 volunteers with expertise in image and video processing. Participants are asked to select the most visually-appealing and semantically-faithful video from those generated by our proposed method and the state-of-the-art approaches, focusing on video quality, consistency with the text, and motion quality. Specifically, each participant has 15 questions. As shown in Table 3, I4VGen performs favorably against the other methods.
@@
Methods Time
AnimateDiff [25] 21.73s
AnimateDiff + FreeInit [77] 113.67s
@@
Figure 5: Intermediate results visualization.
Figure 6: Adaptation on SparseCtrl.
@@ 5.4 Ablation study
On generation-selection strategy. We adopt a generation-selection strategy to create visually-realistic and semantically-faithful anchor images, which serve as a foundation for providing appearance information to enhance subsequent video generation performance. As shown in Table 2, highlighted in orange, compared to randomly synthesizing a single anchor image, the generation-selection strategy significantly improves the quality of the generated videos in terms of frame-wise quality and consistency with the text. Fig. 6 provides a visualization of the candidate images, where the reward-based selection mechanism eliminates unsatisfactory images, leading to better results.
@@
On video regeneration. Fig. 6 visualizes the intermediate results, demonstrating that the video regeneration process is essential for refining appearance details and optimizing motion patterns.
@@ 5.5 More applications
Figure 7: Adaptation on real image.
Adaptation on image-to-video diffusion models. I4VGen can be seamlessly integrated into existing image-to-video diffusion models by replacing the anchor images with the provided images, thereby enhancing the overall video quality. As shown in Fig. 6, integrating I4VGen into SparseCtrl [24] significantly improves the quality of the generated videos in terms of temporal consistency and appearance fidelity.
@@ 6 Conclusion
The paper introduces I4VGen, a training-free and plug-and-play video diffusion inference framework enhancing text-to-video generation through advanced image techniques. It decomposes the video generation process into anchor image synthesis and anchor image-guided video synthesis, utilizing a generation-selection pipeline for the realistic anchor image and noise-invariant video score distillation sampling for animation. I4VGen effectively alleviates non-zero terminal signal-to-noise ratio issues and demonstrates improved visual realism and textual fidelity when integrated with existing video diffusion models.
@@
Limitation and discussion. I4VGen improves the video diffusion model but requires more inference cost. As discussed in Sec. 5.3, the inference time of I4VGen is over double the baseline. Enhancing inference efficiency remains a future goal, with distillation techniques as a potential approach. Additionally, although our method and FreeInit are orthogonal, integrating both by replacing video regeneration with FreeInit fails to produce notable benefits.

### v2

@@ 3.2 Anchor image-augmented video synthesis
  • Instead of resampling the Gaussian noise at each iteration as in traditional SDS, we maintain a constant noise across the optimization, enhancing convergence speed.
  • Optimization is confined to the initial stages of the denoising process, where noise levels are higher, focusing on dynamic information distillation.
  • We implement a coarse-to-fine optimization strategy, evolving from high to low noise levels, specifically from timestep $T$ to $\tau_{\text{NI-VSDS}}$, where $T>\tau_{\text{NI-VSDS}}>0$. This approach stabilizes the optimization trajectory and yields superior motion quality.
Input: T2V diffusion model $\epsilon_{\theta}(\cdot)$, text prompt $\mathbf{y}$, static video $\hat{\mathbf{v}}$, timestep $\tau_{\text{NI-VSDS}}$.
Output: Dynamic video.
@@ 3       $\hat{\mathbf{v}}_{t}\leftarrow\mathsf{AddNoise}\left(\hat{\mathbf{v}},\epsilon,t\right)$
Algorithm 1 NI-VSDS
The implementation of noise-invariant video score distillation sampling (NI-VSDS) algorithm is detailed in Algorithm 1, which outlines the process of converting a static video into a dynamic video using the defined NI-VSDS loss. Notably, we only perform a single update from timestep $T$ to $\tau_{\text{NI-VSDS}}$, requiring fewer than 50 iterations, this is a significant reduction compared to the thousands of iterations typically required for text-to-3D synthesis in SDS. $\alpha$ is a scalar that defines the step size of the gradient update. We empirically set $\tau_{\text{NI-VSDS}}=\mathsf{Int}(T\times p_{\text{NI-VSDS}})$.
AnimateDiff 87.11% 95.22% 95.99% 93.12% 74.89% 56.07% 64.29% 83.69%
+ FreeInit 90.45% 96.57% 96.89% 95.66% 70.17% 59.25% 63.51% 87.55%
+ I4VGen 95.17% 97.73% 98.51% 96.45% 57.72% 64.68% 66.18% 92.59%
LaVie 91.65% 96.30% 98.03% 95.73% 71.94% 59.64% 65.13% 91.25%
+ FreeInit 92.32% 96.35% 98.06% 95.83% 71.11% 59.41% 63.89% 89.13%
+ I4VGen 94.12% 96.90% 98.55% 96.37% 70.55% 60.88% 66.55% 92.26%
AnimateDiff 22.61% 90.40% 81.73% 31.55% 45.61% 24.40% 24.49% 25.71%
+ FreeInit 26.92% 93.00% 86.39% 30.71% 44.61% 23.98% 25.03% 25.61%
+ I4VGen 57.22% 95.80% 91.98% 45.20% 54.67% 25.07% 26.11% 28.01%
LaVie 24.02% 94.80% 83.64% 26.27% 52.89% 23.67% 24.94% 27.25%
+ FreeInit 22.59% 94.20% 84.34% 27.46% 52.70% 23.61% 24.85% 26.89%
+ I4VGen 32.77% 96.20% 88.59% 33.81% 55.64% 24.35% 25.62% 27.68%
Video regeneration. After animating the static video, we further enhance the appearance detail quality of the video through a diffusion-denoising process. This stage is not affected by the aforementioned training-inference gap, thereby achieving more refined generation results.
Notably, we can flexibly add noise up to any timestep $\tau_{\text{re}}$, calculated as $\tau_{\text{re}}=\mathsf{Int}(T\times p_{\text{re}})$, followed by the corresponding denoising process. This strategy not only preserves the fine appearance textures but also reduces the required denoising steps, thus streamlining the video synthesis process and elevating the overall quality of the resulting video.
@@ 4.1 Experimental settings
Implementation details. I4VGen a novel video diffusion inference pipeline that leverages advanced image techniques to enhance pre-trained text-to-video diffusion models without requiring additional training, and can be seamlessly integrated into existing text-to-video diffusion models. To ascertain the efficacy and adaptability of I4VGen, we apply it to two well-regarded text-to-video diffusion models: AnimateDiff (Guo et al., 2024b) and LaVie (Wang et al., 2023e).
@@
AnimateDiff 87.11% 95.22% 95.99% 93.12% 74.89% 56.07% 64.29% 83.69%
+ I4VGen (w/o gen.-sel.) 94.89% 97.80% 98.28% 96.99% 55.91% 62.23% 64.18% 90.95%
+ I4VGen (w/o NI-VSDS) 96.47% 98.82% 98.99% 97.56% 28.24% 65.17% 65.52% 92.66%
+ I4VGen 95.17% 97.73% 98.51% 96.45% 57.72% 64.68% 66.18% 92.59%
@@
AnimateDiff 22.61% 90.40% 81.73% 31.55% 45.61% 24.40% 24.49% 25.71%
+ I4VGen (w/o gen.-sel.) 40.68% 94.40% 90.55% 37.79% 53.72% 24.76% 26.03% 26.62%
+ I4VGen (w/o NI-VSDS) 62.84% 94.80% 91.95% 47.57% 55.80% 24.88% 25.72% 27.91%
+ I4VGen 57.22% 95.80% 91.98% 45.20% 54.67% 25.07% 26.11% 28.01%
For AnimateDiff, the mm-sd-v15-v2 motion module [^1] https://github.com/guoyww/AnimateDiff, alongside Stable Diffusion v1.5, is utilized to synthesize 16 consecutive frames at a resolution of $512\times 512$ pixels for evaluation. For LaVie, the base-version [^2] https://github.com/Vchitect/LaVie is employed to generate 16 consecutive frames at $320\times 512$ pixels for evaluation. All other inference details adhere to the original settings described in Guo et al. (2024b) and Wang et al. (2023e), respectively. Notably, both AnimateDiff and LaVie possess inherent text-to-image generation capabilities when excluding the motion module. To avoid introducing additional GPU storage requirements, we leverage their corresponding image versions for text-to-image generation in I4VGen. For AnimteDiff, we empirically set $N=16$, $p_{\text{NI-VSDS}}=0.4$, $\alpha=1$, and $p_{\text{re}}=1$. For LaVie, we empirically set $N=16$, $p_{\text{NI-VSDS}}=0.4$, $\alpha=1$, and $p_{\text{re}}=0.8$. All experiments are conducted on a single NVIDIA V100 GPU (32 GB).
Benchmark. I4VGen is assessed using VBench (Huang et al., 2024), a comprehensive benchmark that evaluates video generation models across 16 disentangled dimensions, which is more authoritative than FVD. These dimensions provide a detailed analysis of generation quality from two overarching perspectives: video quality [^3] Video quality includes 7 evaluation dimensions: Subject Consistency, Background Consistency, Temporal Flickering, Motion Smoothness, Dynamic Degree, Aesthetic Quality, and Imaging Quality. The first 5 evaluate temporal quality, and the last 2 evaluate frame-wise quality., focusing on the perceptual quality of the generated videos, and video-condition consistency [^4] Video-condition consistency includes 9 evaluation dimensions: Object Class, Multiple Objects, Human Action, Color, Spatial Relationship, Scene, Appearance Style, Temporal Style, Overall Consistency. The first 6 evaluate semantics, the 7 and 8-th evaluate style, and the 9-th evaluates overall consistency., assessing how well the generated videos align with the provided conditions.
4.2 Qualitative comparison
Fig. 4 presents a comparative analysis of our results against state-of-the-art counterparts using identical text prompts and random seeds. I4VGen excels in enhancing both the temporal consistency and the frame-wise quality, alongside superior alignment with the text prompts. For instance, in the case of “playing guitar", AnimateDiff suffers from poor video quality, and FreeInit encounters an incomplete guitar in the middle of the video. In contrast, our method effectively addresses these issues, maintaining stable temporal consistency. Furthermore, while baseline methods struggle with accurate synthesis of all text-described components, e.g., “NYC Times Square", I4VGen generates videos that are visually realistic and closely aligned with the text prompts by utilizing anchor images obtained by the generation-selection strategy.
@@ 4.3 Quantitative comparison
Objective evaluation. Following the protocols established by VBench, we evaluate I4VGen in terms of both video quality and video-text consistency. As detailed in Table 1, I4VGen outperforms all other approaches in temporal quality (higher background and subject consistency, less flickering, and better smoothness), frame-wise quality (higher aesthetic and imaging quality), and video-text consistency (greater semantics, style, and overall consistency). Although counterparts occasionally produce videos with more dynamic motion, they are often linked to inappropriate or excessive movements. I4VGen strikes a more effective balance between motion intensity and overall video quality, which is further verified in the user study.
Table 3: User study.
Method Video Quality Vid.-Cond. Consistency Overall score
AnimateDiff 6.00% 10.67% 6.33%
+ FreeInit 27.67% 15.67% 25.00%
+ I4VGen 66.33% 73.67% 68.67%
LaVie 27.67% 21.33% 22.33%
+ FreeInit 22.67% 18.33% 19.67%
+ I4VGen 49.67% 60.33% 58.00%
User study. We conduct a subjective user study involving 20 volunteers with expertise in image and video processing, with each participant answering 15 questions. Specifically, participants are asked to select the video with the highest quality across three dimensions: video quality, video-condition consistency, and overall score. As shown in Table 3, our approach outperforms the other methods favorably.
@@
Method Time
AnimateDiff 21.73s
AnimateDiff + FreeInit 113.67s
@@
Figure 5: Intermediate results visualization. We provide visualizations of the candidate images with reward scores, the dynamic video, and the corresponding generated video.
@@ 4.4 Ablation study
On generation-selection strategy. We adopt a generation-selection strategy to create visually-realistic and semantically-faithful anchor images, which serve as a foundation for providing appearance information to enhance subsequent video generation performance. As shown in Table 2, highlighted in orange, compared to randomly synthesizing a single anchor image, the generation-selection strategy significantly improves the quality of the generated videos in terms of frame-wise quality and consistency with the text. Fig. 5 provides a visualization of the candidate images, where the reward-based selection strategy eliminates unsatisfactory images, leading to better results.
@@
Figure 6: Adaptation on SparseCtrl. I4VGen can be seamlessly integrated into SparseCtrl by replacing the anchor image with the provided image, leading to improved results.
On video regeneration. Fig. 5 visualizes the intermediate results, demonstrating that the video regeneration process is essential for refining appearance details.
Figure 7: Adaptation on real image.
@@ 4.5 More applications
Adaptation on image-to-video diffusion models. I4VGen can be seamlessly integrated into existing image-to-video diffusion models by replacing the anchor images with the provided images, thereby enhancing the overall video quality. As shown in Fig. 6, integrating I4VGen into SparseCtrl (Guo et al., 2023) significantly improves the quality of the generated videos in terms of temporal consistency and appearance fidelity.
@@ 5 Conclusion
The paper introduces I4VGen, a novel video diffusion inference pipeline to leverage advanced image techniques to enhance pre-trained text-to-video diffusion models, which requires no additional learnable parameters and training costs. I4VGen decomposes the text-to-video inference process into anchor image synthesis and anchor image-augmented video synthesis. Correspondingly, a simple yet effective generation-selection strategy is applied to produce a high-quality anchor image, and an innovative noise-invariant video score distillation sampling (NI-VSDS) is designed to animate the image, followed by a video regeneration process to enhance the final output. I4VGen effectively alleviates non-zero terminal signal-to-noise ratio issues and demonstrates improved visual realism and textual fidelity when integrated with existing video diffusion models.
@@
Limitation and discussion. I4VGen improves the video diffusion model but requires more inference cost. As discussed in Sec. 4.3, the inference time of I4VGen is over double the baseline. Enhancing inference efficiency remains a future goal, with distillation techniques as a potential approach. Furthermore, removing the generation-selection strategy can reduce inference costs to some extent. As shown in Table 2, our method still significantly outperforms the baseline under this setting. Additionally, although our method and FreeInit are orthogonal, integrating both by replacing video regeneration with FreeInit fails to produce notable benefits.
Appendix A Related Work
Video Generative Models. The domain of video generation has seen significant advancements through the use of Generative Adversarial Networks (GANs) (Vondrick et al., 2016; Saito et al., 2017; Tulyakov et al., 2018; Wang et al., 2020; Saito et al., 2020; Tian et al., 2021; Fox et al., 2021; Yu et al., 2022; Skorokhodov et al., 2022; Brooks et al., 2022; Shen et al., 2023; Wang et al., 2023f), Variational Autoencoders (VAEs) (Mittal et al., 2017; Li et al., 2018; He et al., 2018), and Autoregressive models (ARs) (Yan et al., 2021; Ge et al., 2022; Wu et al., 2022; Hong et al., 2023; Villegas et al., 2023; Fu et al., 2023; Yoo et al., 2023; Yu et al., 2023b). Despite these developments, synthesizing videos from text prompts remains challenging due to the complexities of modeling spatio-temporal dynamics. Recent innovations driven by the successes of diffusion models (Ho et al., 2020; Dhariwal & Nichol, 2021; Song et al., 2021), which have been applied effectively in image generation (Rombach et al., 2022; Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Gu et al., 2023; Balaji et al., 2022; Xue et al., 2023; Meng et al., 2022; Guo et al., 2024a) and audio synthesis (Kong et al., 2021; Chen et al., 2021; Popov et al., 2021; Leng et al., 2022; Liu et al., 2022), and underscore the emergence of substantial headway (Ho et al., 2022b; a; He et al., 2022; Singer et al., 2023; Blattmann et al., 2023b; Yu et al., 2023c; Ruan et al., 2023; Wu et al., 2023a; Chen et al., 2023a; b; Esser et al., 2023; Ge et al., 2023; Chen et al., 2024a; Geyer et al., 2024; Ma et al., 2023; Wang et al., 2023e; Zhang et al., 2023a; b; Hu et al., 2023; Wang et al., 2023d; Feng et al., 2023; Guo et al., 2024b; Girdhar et al., 2023; Blattmann et al., 2023a; Gupta et al., 2023; Wang et al., 2023b; c; Luo et al., 2023) in research endeavors devoted to video synthesis from text input.
The foundational contributions of the Video Diffusion Model (VDM) (Ho et al., 2022b) represents a milestone in leveraging diffusion models for video generation by adapting the 2D U-Net architecture used in image generation to a 3D U-Net capable of temporal modeling. Successive researches, such as Make-A-Video (Singer et al., 2023) and Imagen Video (Ho et al., 2022a), expand video generation capabilities significantly. To enhance efficiency, subsequent models have transitioned the diffusion process from pixel to latent space (He et al., 2022; Zhou et al., 2022; Wang et al., 2023b; Blattmann et al., 2023b; a; Guo et al., 2024b; Wang et al., 2023e), paralleling advancements in latent diffusion for images (Rombach et al., 2022).
However, the direct generation of videos from text prompts remains intrinsically challenging. Recent approaches (Blattmann et al., 2023a; Zhang et al., 2023b; Girdhar et al., 2023; Chen et al., 2023a; 2024a; Li et al., 2023; Hu et al., 2023; Yu et al., 2023a; Ren et al., 2024) have employed text-to-image synthesis as an intermediary step, enhancing overall performance. Despite these advancements, these methods still face the challenge of high computational training costs. In this study, we explore a novel training-free methodology aimed at bridging the existing gap in the field.
In addition, (Chen et al., 2024b) (contemporary researches) introduces additional operations in the attention layer, i.e., cross-frame self-attention control, to enhance the video model. However, this necessitates modifications to the model architecture, whereas our method does not.
Unfortunately, most existing diffusion models exhibit a disparity between the corrupted image during training and the pure Gaussian noise during inference. Commencing denoising from pure Gaussian noise in the inference phase deviates from the training process, potentially introducing signal-leak bias. For image diffusion models, (Guttenberg, ; Lin et al., 2024; Li et al., 2024) point out flaws in common diffusion noise schedules and sample steps, and propose to fine-tune the diffusion model to mitigate or eliminate the signal-leak bias during training, leading to improved results. (Everaert et al., 2024) attempts to exploit signal-leak bias to achieve more control over the generated images. For video diffusion models, this issue becomes more pronounced. (Wu et al., 2023b; Ma et al., 2023) invert the retrieved video or generated low-quality to construct initial noise to alleviate the problem of signal-leak, improving inference quality. However, they suffer from limited diversity and cumbersome inference. At the same time, first-round inference of FreeInit (Wu et al., 2023b) still exhibits a training-inference gap.
Figure 9: Impact of hyperparameters. For different texts, the optimal parameter settings are different, and the sensitivity to parameters also varies. However, they all significantly outperform the baseline. In this paper, we provide an empirical setting that is mild for most cases, serving as a performance lower bound for I4VGen. I4VGen supports fine-tuning parameters on a per-example, achieving higher-quality videos.
Figure 10: Failure cases. I4VGen is designed to fully unleash the potential of existing video diffusion models, but it still cannot synthesize high-quality videos that are out of the distribution. For example, the text marked in red.
Appendix B Experiments
B.1 Qualitative comparison
We provide more visualization results in Fig. 8, it can be seen that our method generates more semantically plausible and photo-realistic results than its counterparts. We provide the videos shown in the main paper and appendix in mp4 format in the Supplementary material.
B.2 Quantitative comparison
On hyperparameters. I4VGen is a training-free method that improves video generation performance by correcting the inference process. It is obvious that I4VGen is also a case-wise method, where different cases correspond to different optimal hyperparameters. In this paper, we provide an empirical setting that is mild for most instances, serving as a performance lower bound for I4VGen, and facilitating large-scale quantitative comparisons. Furthermore, we also provide a visualization of the impact of hyperparameters in Fig. 9, which shows that carefully tuned hyperparameters can achieve higher-quality videos.
B.3 Failure cases and discussions
We provide the failure cases in Fig. 10, I4VGen is designed to fully unleash the potential of existing video diffusion models, but it still fails to synthesize high-quality videos that are out of the distribution.
Appendix C Code
We also provide the code for I4VGen in the Supplementary material.